OpenStudy (anonymous):

anjo works 5 mins. faster than aira in cleaning the room. if they work together, they can finish the job in 6 mins. How long can each of them clean the room working alone?

OpenStudy (saifoo.khan):

\[\Large \frac{1}{x}+\frac{1}{x+5}=\frac16\]Solve for x.

OpenStudy (anonymous):

then ?

OpenStudy (saifoo.khan):

then, take the LCM and solve.

OpenStudy (anonymous):

whats the next steps ? please ???

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

let anjo can finish in x min so aira can finish in x+5 min in 1 min anjo can do 1/x part of work in 1 min aira can do 1/(x+5) part of work togather in 1 min they can do 1/x + 1/(x+5) part of work they finished togather in 6 min so 6[1/x + 1/(x+5)] = 1 (6/x)+(6/(x+5))=1 6(x+5)+6(x)=(x)(x+5) 6x+30+6x=x^2+5x x^2-7x-30=0 (x-10)(x+3)=0 x=-3 and x=10 time cannot be negative so x=10 x+5= 10+5=15